40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 9 00:14:02 2012, Eric, Summary, Locking, YARM Locking and Calibration cal_swept_sine3_tfmagcal_swept_sine3_tfphcal_swept_sine3_cohsensing_func_model.png
    Reply  Thu Aug 9 00:35:53 2012, Jenne, Summary, Locking, YARM Locking and Calibration 
       Reply  Thu Aug 9 10:09:32 2012, Eric, Summary, Locking, YARM Locking and Calibration cal_swept_sine3_magnitude.pdfcal_swept_sine3_phase.pdfcal_swept_sine3_coherence.pdf
Message ID: 7128     Entry time: Thu Aug 9 00:14:02 2012     Reply to this: 7130
Author: Eric 
Type: Summary 
Category: Locking 
Subject: YARM Locking and Calibration 

Today I spent time locking the YARM in order to calibrate the arm cavity. Here's what I did:

1. Misalign all optics other than the beam splitter, ITMY, ETMY and PZT2

2. Restore BS, ITMY, ETMY, and PZT2

3. Open Dataviewer and run /users/Templates/JenneLockingDataviewer/Yarm.xml from the Restore Settings. This opens the signals C1:LSC-POY11_I_ERR (the Pound-Drever-Hall error signal for this measurement) and C1:LSC-TRY_OUT (the light transmitted through ETMY) in the plot window.

4. Adjust ITMY and ETMY pitch and yaw using the video screens looking at AS and ETMYT as a first, rough guide. It can be helpful at first to increase the gain on the YARM servo filter module in the C1LSC control screen to about 0.3 and decrease it back down to 0.1 as the beam becomes better aligned. You know when to decrease this gain when fuzzy, small oscillations appear on the C1:LSC-TRY_OUT signal. If the mode cleaner is locked you should see a bright spot on the AS camera.

5. Tinker with pitch and yaw while looking at the AS screen until you see a reasonably good circular spot without other fringes extending from a bright center.

6. The overall goal is to maximize C1:LSC-TRY_OUT because the power transmitted through EMTY is proportional to the power within the cavity. A decent target value is 0.85 and today I was able to get it to just over 0.80 at best. At first there will probably be small spikes in C1:LSC-TRY_OUT. You want to adjust pitch and yaw until the deviation in the signal from zero is no longer just a spike, but a sustained, flat signal above zero. By this time there should be light showing up on the ETMYT camera as well.

7. Once that happens, continue to successively adjust ITMY and ETMY doing the pitch adjustments on both first, and then the yaw adjustments, or vice versa. You can also tweak the PZT2 pitch and yaw. Once you've got C1:LSC-TRY_OUT as large as possible, you've locked the cavity.

I saved the pitch and yaw settings I ended up with for ITMY, ETMY, BS and PZT2 in the IFO_ALIGN screen. Before the end of the day I think Jenne restored the rest of the previously misaligned optics because they were restored when I got back from dinner.

 

 

I also worked on calibrating the YARM. I opened up DTT using C1:LSC-POY11_I_ERR as the measurement channel and C1:SUS-ITMY_LSC_EXC as the excitation channel. I ran a logarithmic swept sine response measurement with 100 points and an amplitude of 25. The mode cleaner kept losing its lock all day, and if this happened while making this measurement I tried to pause the sweep as quickly as possible. I analyzed the the transfer function and the coherence function that the sweep produced, and thought that some of the odd behavior was due to losing the lock and getting back to a slightly different locked state when resuming the measurement. The measured transfer function and coherence plots are attached below. Both the transfer function and the coherence look good above roughly 30 Hz, but do not look correct at low frequencies. There's also a roll-off in the measured transfer function around 200 Hz, while in the model the magnitude of the transfer function drops only after the corner frequency of the cavity, around several kHz. I have attached a plot of the roughly analogous transfer function from the DARM control loop model (the gains are very large due to the large arm cavity gain and the ADC conversion factor of 2^16/(20 V) ). The measured and the modeled transfer functions are slightly different in that the model does not include the individual mirrors, while the excitation was imposed on ITMY for the measurement.

 

The next steps are to figure out what's happening in DTT with the transfer function and coherence at low frequencies, and to understand the differences between the model and the measurement.

Attachment 1: cal_swept_sine3_tfmag  90 kB
Attachment 2: cal_swept_sine3_tfph  89 kB
Attachment 3: cal_swept_sine3_coh  72 kB
Attachment 4: sensing_func_model.png  26 kB  | Hide | Hide all
sensing_func_model.png
ELOG V3.1.3-